SHOW WORK!

The scores on two standardized tests are normally distributed.
The first test had a mean of 56 and a standard deviation of 6.
The second test had a mean of 76 and a standard deviation of 6.
What score would you need on the second test to equal a score of 70 on the first test?
Give answer to the nearest whole number.
Answer #1

Calculating the z score for the first test:

Z = \frac{X - \mu}{\sigma} = \frac{70-56}{6} = \frac{7}{3}=2.3333

Now we have to find Y such that:

Z = \frac{Y - \mu}{\sigma} = \frac{Y-76}{6} = \frac{7}{3}

\rightarrow \frac{Y-76}{6} = \frac{7}{3}

\rightarrow Y = 90

Hence 90 score would be needed on the second test to equal a score of 70 on the first test.

Let me know in the comments if anything is not clear. I will reply ASAP! Please do upvote if satisfied!